The region b/w  x=0 &x=s is filled with uniform steady magnetic field b(k cap).aparticle of mass m,+ve charge q and velocity v(i cap)travels along x-axis and enters the region of the magnetic field


neglect the gravity throught the question


(a) find the value of s if the partical emerges from the region of magnetic field with final velocity at an angle of 30(degree) to its initial velocity.


(b)find the final velocity of the particle and the time spent by it in magnetic field,if the magnetic field now extends upto 2.1s.                            

Grade:12

1 Answers

Nagaraju Kandru
14 years ago

sd

 

 (a)  the deviation of a charged particle in a uniform magnetic field is sinθ = s/r

                  where 's' is the length of the magnetic field.

       (for derivation plzz see ARIHANT: electricity  and magnetism by D.C. PANDEY)

          here deviation is 30o . and sin30o = 0.5= s/r

                         =>   s = r/2 where r is  the radius of circular arc traced by the particle

                           r = mv/bq.

 

(b)  plzz make the things clear like velocity(magnitude, direction or both) and 2.1s .